# From affine algebraic racks to Leibniz algebras

### v1

@@ 4.3. Properties of corack algebras
For example, let $Q=\operatorname{Spec}(A)$ be an affine algebraic rack. Then for all $k$-algebras $R$ and $R$-points $x\in Q(R)$ (thought of as $k$-algebra homomorphisms $x^{\sharp}\colon A\to R$), the automorphism $L_{x}\colon Q(R)\xrightarrow{\sim}Q(R)$ from (2.4) induces an automorphism
@@
(4.3) $r\otimes f$ $\mapsto(r\otimes 1)\cdot((x^{\sharp}\otimes\operatorname{id}_{A})\circ\nabla)(f).$
4.3.1.
As another example, we show that corack algebras are not coassociative or cocommutative except in trivial cases. This stands in stark contrast to Hopf algebras of affine algebraic groups. Mirroring that setting, let us call a corack algebra $A$ coassociative if $(\mathord{\operatorname{id}}\otimes\nabla)\circ\nabla=(\nabla\otimes\operatorname{id})\circ\nabla$ and cocommutative if $\tau\circ\nabla=\nabla$, where $\tau\colon A\otimes A\to A\otimes A$ is the transposition $a\otimes b\mapsto b\otimes a$.
@@
for all $x,y,z\in Q$. Since $L_{x\triangleright y}$ is bijective (see the discussion around (2.4)), the claim follows. ∎
@@
Under these conditions, $\nabla=1\otimes\operatorname{id}$ as in Example 4.4.
@@
Certainly, if $Q$ is a singleton (that is, if condition 3 holds), then $\triangleright$ is commutative and $\triangleright=\triangleright^{-1}$. The converses follow from the fixing and fixedness axioms 4 and 5: if $\triangleright$ is commutative, then
@@
for all $x\in Q$, and if $\triangleright=\triangleright^{-1}$, then
$x=e\triangleright x=e\triangleright^{-1}x=L_{x}^{-1}(e)=e$
@@
If $Q=\operatorname{Spec}(A)$ is an affine algebraic rack, then $Z(Q)$ is a closed algebraic left ideal of $Q$.
@@
$1\otimes f=L_{x}^{\sharp}(1\otimes f)=(x^{\sharp}\otimes\operatorname{id}_{A})(\nabla f)$
@@
Let $G$ be an affine algebraic group. Then $Z(\operatorname{Conj}(G))=Z(G)$ is a closed algebraic normal subgroup of $G$ and, hence, a closed algebraic left ideal of $\operatorname{Conj}(G)$.
@@ 5.2. Tangent spaces
We recall certain facts about tangent spaces and tangent maps of affine schemes; see [milne]*App. A.h and [borel]*AG.16 for details. Henceforth, let $Q=\operatorname{Spec}(A)$ be an affine algebraic rack.
5.2.1.
@@
The Leibniz algebra of $Q$ is the vector space
@@
equipped with the Leibniz bracket $\cdot,\cdot]\colon\mathfrak{q}\times\mathfrak{q}\to\mathfrak{q}$ defined by
@@
Writing $\nabla f=\sum f_{(1)}\otimes f_{(2)}$, we thus have
$D,E](f)=\sum D(f_{(1)})E(f_{(2)})$
for all $f\in A$. Note that $\cdot$$\cdot$] is $k$-bilinear.
5.2.2.
@@
$\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(V)),k)$ $\xrightarrow{\sim}T_{\operatorname{id}}\mathord{\operatorname{GL}}(V)$ $\mathfrak{gl}(V)=\operatorname{End}(V)$ $\xrightarrow{\sim}T_{\operatorname{id}}\mathord{\operatorname{GL}}(V)$
(5.1) $F$ $\mapsto\varepsilon_{\mathcal{O}(\operatorname{GL}(V))}+\delta F,$ $f$ $\mapsto\operatorname{id}_{k[\delta]\otimes_{k}V}+\delta f.$
In the isomorphism on the left, we identify $T_{\operatorname{id}}\mathord{\operatorname{GL}}(V)$ as the vector space of $k$-algebra homomorphisms $\mathcal{O}(\operatorname{GL}(V))\to k[\delta]$ whose composite with $k[\delta]\to k$ is the counit $\varepsilon_{\mathcal{O}(\operatorname{GL}(V))}$ of the Hopf algebra $\mathcal{O}(\operatorname{GL}(V))$. On the right, we dually identify $T_{\operatorname{id}}\mathord{\operatorname{GL}}(V)$ as the subgroup
$T_{\operatorname{id}}\mathord{\operatorname{GL}}(V)=\ker(\operatorname{GL}(V)(k[\delta])\to\operatorname{GL}(V)(k))\leq\operatorname{GL}(V)(k[\delta]).$
@@
In particular, to pass from $\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(\mathfrak{q})),k)$ to $\mathfrak{gl}(\mathfrak{q})$ during the proof of Theorem 1.1 using (5.1), we will need to pass between these two equivalent interpretations of $T_{\operatorname{id}}\mathord{\operatorname{GL}}(\mathfrak{q})$. See Section 5.3.1 for an overview of this argument.
@@ 5.2.4. Tangent maps
Given a morphism of schemes $\varphi\colon X\to Y$ at $x\in X$, let $\varphi^{\sharp}\colon\mathcal{O}(Y)\to\mathcal{O}(X)$ be the pullback $k$-algebra homomorphism. Identify $T_{e}X$ and $T_{\varphi(x)}Y$ with the respective spaces of $k$-derivations. Then the derivative, tangent map, or differential of $\varphi$ at $x$ is the $k$-linear map $(d\varphi)_{x}\colon T_{x}X\to T_{\varphi(x)}Y$ defined by
@@ 5.3. Adjoint representations
Recall from (2.4) that every point $x\in Q$ induces an automorphism of algebraic racks $L_{x}\colon Q\xrightarrow{\sim}Q$. It follows that the derivative $(dL_{x})_{e}$ is a $k$-linear automorphism of $\mathfrak{q}$, allowing us to define the following.
@@
5.3.1.
For the purposes of proving Theorem 1.1, Definition 5.3 is problematic because it defines the codomain of $\operatorname{ad}$ to be $\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(\mathfrak{q})),k)$ rather than $\mathfrak{gl}(\mathfrak{q})$. To find the element of $\mathfrak{gl}(\mathfrak{q})$ corresponding to $\operatorname{ad}_{D}\in\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(\mathfrak{q})),k)$ for each $k$-derivation $D\in\mathfrak{q}$, we employ the isomorphisms in Proposition 5.2 and (5.1). Section 6.1 is dedicated to this argument.
Here is an overview of the argument in Section 6.1. Given $D\in\mathfrak{q}$, we lift $D$ to a $k$-algebra homomorphism $\partial D\colon A\to k[\delta]$ as in Proposition 5.2. Viewing $\partial D$ as a $k[\delta]$-point of $Q$ gives us a $k[\delta]$-point $\operatorname{Ad}_{\partial D}\in\operatorname{GL}(k[\delta]\otimes_{k}\mathfrak{q})$ of $\operatorname{GL}(\mathfrak{q})$. Then the second isomorphism in (5.1) states that for all $k$-derivations $E\in\mathfrak{q}$, the coefficient of $\delta$ in $\operatorname{Ad}_{\partial D}(1\otimes E)\in k[\delta]\otimes_{k}\mathfrak{q}\cong\mathfrak{q}[\delta]/(\delta^{2})$ is precisely the element of $\mathfrak{q}$ corresponding to $\operatorname{ad}_{D}(E)$. We show that this coefficient is the convolution $[D,E]$.
@@ 6.1. Convolution versus $\operatorname{ad}$
6.1.1.
In the following, fix arbitrary $k$-derivations $D,E\in\mathfrak{q}$. Our proofs of Propositions 6.2 and 7.9 will use the $k$-linear map $\psi_{D}\colon A\to A$ defined by
(6.1) $\psi_{D}\coloneq(D\otimes\operatorname{id})\circ\nabla.$
@@
Via the identification $k[\delta]\otimes_{k}A\xrightarrow{\sim}A[\delta],$ we obtain from (4.3) an automorphism $L_{\partial D}^{\sharp}$ of $A[\delta]$ defined by
$L_{\partial D}^{\sharp}(\delta^{n}f)=\delta^{n}(\partial D\otimes\operatorname{id}_{A})(\nabla f)$
for all $f\in A$. The following is an alternative formula for $L_{\partial D}^{\sharp}$.
@@
$L_{\partial D}^{\sharp}(\delta^{n}f)=\delta^{n}(f+\delta\psi_{D}(f)).$
@@
$=\delta^{n}\sum\varepsilon(f_{(1)})f_{(2)}+\delta\sum D(f_{(1)})f_{(2)}$
@@
where in the last equality we have used the co-fixing identity 4. ∎
6.1.3.
Similarly, since $\partial D$ is a $k[\delta]$-point of $Q$ and $\operatorname{Ad}\colon Q\to\operatorname{GL}(\mathfrak{q})$ is a morphism of schemes, we obtain a $k[\delta]$-point of $\operatorname{GL}(\mathfrak{q})$
$\operatorname{Ad}_{\partial D}\in\operatorname{GL}(\mathfrak{q})(k[\delta])=\operatorname{GL}(k[\delta]\otimes_{k}\mathfrak{q}).$
In particular, under the identification $k[\delta]\otimes_{k}A\xrightarrow{\sim}A[\delta]$, the definition of the derivative shows that
(6.3) $\operatorname{Ad}_{\partial D}(1\otimes E)=E\circ L_{\partial D}^{\sharp}\colon A[\delta]\to k[\delta]$
for all $k$-derivations $E\in\mathfrak{q}$. We use this to show the following.
@@
When considered as a map $\operatorname{ad}\colon\mathfrak{q}\to\mathfrak{gl}(\mathfrak{q})$, the adjoint representation of $\mathfrak{q}$ is given by the left argument of the Leibniz bracket $D\mapsto[D,\cdot]$. That is, $\operatorname{ad}_{D}(E)=[D,E]$ for all $k$-derivations $D,E\in\mathfrak{q}$. In particular, $\mathfrak{q}$ is closed under $\cdot$$\cdot$].
@@
If $G$ is an affine algebraic group and $Q=\operatorname{Conj}(G)$ is its conjugation quandle scheme, then $(\mathfrak{q},[\cdot,\cdot])$ is the Lie algebra of $G$. To see this, note that $Q$ and $G$ have the same underlying scheme and identity $e$, so they also have the same tangent space $\mathfrak{q}$ at $e$. By the definition of $\operatorname{Conj}(G)$, the adjoint representations $\operatorname{Ad}$ of $Q$ and $\operatorname{ad}$ of $\mathfrak{q}$ are therefore the same as those of $G$ and its Lie algebra (see, for example, [milne]*Rem. 10.24). Hence, the claim follows from Proposition 6.2.444Alternatively, one can use Example 4.5 and the Hopf algebra axioms to directly (if tediously) verify that the Leibniz bracket $D,E]=(D\otimes E)\circ\nabla$ agrees with the usual Lie bracket $(D\otimes E-E\otimes D)\circ\Delta$ on $\mathfrak{q}=\operatorname{Lie}(G)$ (see [waterhouse]*Sec. 12.2); we leave the details to the reader. As a hint, first deduce the identity $\varepsilon\circ S=\varepsilon$ from the axiom $\mu\circ(S\otimes\operatorname{id})\circ\Delta=\eta\circ\varepsilon$ by post-composing with $\varepsilon$, and then deduce from these two equalities that $D\circ S=-D$.

### v2

@@ 6.3. Properties of corack algebras
For example, let $Q=\operatorname{Spec}(A)$ be an affine algebraic rack. Then for all $k$-algebras $R$ and $R$-points $x\in Q(R)$ (thought of as $k$-algebra homomorphisms $x^{\sharp}\colon A\to R$), the automorphisms $L_{x}^{\pm 1}\colon Q(R)\xrightarrow{\sim}Q(R)$ from (3.4) induce automorphisms
@@
(6.3) $r\otimes f$ $\mapsto(r\otimes 1)\cdot((x^{\sharp}\otimes\mathrm{id}_{A})\circ\nabla)(f)$
and
$(L_{x}^{-1})^{\sharp}\colon A\otimes_{k}R$ $\xrightarrow{\sim}A\otimes_{k}R$
(6.4) $f\otimes r$ $\mapsto(1\otimes r)\cdot((\mathrm{id}_{A}\otimes x^{\sharp})\circ\nabla^{-1})(f).$
6.3.1.
As another example, we show that corack algebras are not coassociative or cocommutative except in trivial cases. This stands in stark contrast to Hopf algebras of affine algebraic groups. Mirroring that setting, let us call a corack algebra $A$ coassociative if $(\mathord{\mathrm{id}}\otimes\nabla)\circ\nabla=(\nabla\otimes\mathrm{id})\circ\nabla$ and cocommutative if $\tau\circ\nabla=\nabla$, where $\tau$ denotes the transposition $a\otimes b\mapsto b\otimes a$.
@@
for all $x,y,z\in Q$. Since $L_{x\triangleright y}$ is invertible (see the discussion around (3.4)), the claim follows. ∎
@@
Under these conditions, $\nabla=1\otimes\mathrm{id}$ as in Example 6.4.
@@
Certainly, if $Q$ is a singleton (that is, if condition 3 holds), then $\triangleright$ is commutative and $\triangleright=\triangleright^{-1}$. The converses follow from the fixing and fixedness axioms 4 and 5. Namely, if $\triangleright$ is commutative, then
@@
for all $x\in Q$. On the other hand, if $\triangleright=\triangleright^{-1}$, then the inverse fixedness identity 5 implies that
$x=e\triangleright x=e\triangleright^{-1}x=e$
@@
If $Q=\operatorname{Spec}(A)$ is an affine algebraic rack, then $Z(Q)$ is a closed affine algebraic left ideal of $Q$.
@@
$1\otimes f=L_{x}^{\sharp}(1\otimes f)=(x^{\sharp}\otimes\mathrm{id}_{A})(\nabla f)$
@@
Let $G$ be an affine algebraic group. Then $Z(\operatorname{Conj}(G))=Z(G)$ is a closed algebraic normal subgroup of $G$ and, accordingly, a closed algebraic left ideal of $\operatorname{Conj}(G)$.
@@ 7.2. Tangent spaces
We discuss tangent spaces and tangent maps of affine schemes; see [milne]*App. A.h and [borel]*AG.16 for details. Henceforth, let $Q=\operatorname{Spec}(A)$ be an affine algebraic rack, and recall from Section 6.1 that $Q$ induces a corack algebra structure $(\nabla,\nabla^{-1},\varepsilon)$ on $A$.
7.2.1.
@@
The left Leibniz algebra of $Q$ is the vector space
@@
equipped with the Leibniz bracket $\cdot,\cdot]\colon\mathfrak{q}^{2}\to\mathfrak{q}$ defined by
@@
Similarly, the right Leibniz algebra of $Q$ is the pair $(\mathfrak{q},\{\cdot,\cdot\})$, where
$\{D,E\}\coloneq(D\otimes E)\circ\nabla^{-1}.$
Remark 7.2.
Given a function $f\in A$, write $\nabla f=\sum f_{(1)}\otimes f_{(2)}$. Then
$D,E](f)=\sum D(f_{(1)})E(f_{(2)}),$
and similarly for $\{D,E\}(f)$. This makes it clear that $\cdot$$\cdot$] and $\{\cdot,\cdot\}$ are $k$-bilinear.
7.2.2.
@@
$\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(V)),k)$ $\xrightarrow{\sim}T_{\mathrm{id}}\mathord{\operatorname{GL}}(V)$ $\mathfrak{gl}(V)=\operatorname{End}(V)$ $\xrightarrow{\sim}T_{\mathrm{id}}\mathord{\operatorname{GL}}(V)$
(7.1) $F$ $\mapsto\varepsilon_{\mathcal{O}(\operatorname{GL}(V))}+\delta F,$ $f$ $\mapsto\mathrm{id}_{k[\delta]\otimes_{k}V}+\delta f.$
In the isomorphism on the left, we identify $T_{\mathrm{id}}\mathord{\operatorname{GL}}(V)$ as the vector space of $k$-algebra homomorphisms $\mathcal{O}(\operatorname{GL}(V))\to k[\delta]$ whose composite with $k[\delta]\to k$ is the counit $\varepsilon_{\mathcal{O}(\operatorname{GL}(V))}$ of the Hopf algebra $\mathcal{O}(\operatorname{GL}(V))$. On the right, we dually identify $T_{\mathrm{id}}\mathord{\operatorname{GL}}(V)$ as the subgroup
$T_{\mathrm{id}}\mathord{\operatorname{GL}}(V)=\ker(\operatorname{GL}(V)(k[\delta])\to\operatorname{GL}(V)(k))\leq\operatorname{GL}(V)(k[\delta]).$
@@
In particular, to pass from $\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(\mathfrak{q})),k)$ to $\mathfrak{gl}(\mathfrak{q})$ during the proof of Theorem 1.1 using (7.1), we will need to pass between these two equivalent interpretations of $T_{\mathrm{id}}\mathord{\operatorname{GL}}(\mathfrak{q})$. See Section 7.3.1 for an overview of this argument.
@@ 7.2.4. Tangent maps
Given a morphism of schemes $\varphi\colon X\to Y$ and a point $x\in X$, let $\varphi^{\sharp}\colon\mathcal{O}(Y)\to\mathcal{O}(X)$ be the pullback $k$-algebra homomorphism. Identify $T_{x}X$ and $T_{\varphi(x)}Y$ with the respective spaces of $k$-derivations. Then the derivative, tangent map, or differential of $\varphi$ at $x$ is the $k$-linear map $(d\varphi)_{x}\colon T_{x}X\to T_{\varphi(x)}Y$ defined by
@@ 7.3. Adjoint representations
Recall from (3.4) that every point $x\in Q$ induces automorphisms of algebraic racks $L_{x},L_{x}^{-1}\colon Q\xrightarrow{\sim}Q$. It follows that the tangent maps $(dL_{x})_{e},(dL_{x}^{-1})_{e}$ are $k$-linear automorphisms of $\mathfrak{q}$, allowing us to define the following.
@@
Similarly, define the dual adjoint representations of $Q$ and $\mathfrak{q}$ via
$\operatorname{Ad}^{-1}_{x}\coloneq(dL_{x}^{-1})_{e},\qquad\operatorname{ad}^{-1}\coloneq(d\operatorname{Ad}^{-1})_{e}.$
7.3.1.
For the purposes of proving Theorem 1.1, Definition 7.4 is problematic because it defines the codomain of $\operatorname{ad}^{\pm 1}$ to be $\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(\mathfrak{q})),k)$ rather than $\mathfrak{gl}(\mathfrak{q})$. To find the element of $\mathfrak{gl}(\mathfrak{q})$ corresponding to $\operatorname{ad}_{D}^{\pm 1}\in\operatorname{Der}_{k}(\mathcal{O}(\operatorname{GL}(\mathfrak{q})),k)$ for each $k$-derivation $D\in\mathfrak{q}$, we employ the isomorphisms in Proposition 7.3 and (7.1). Section 8.1 is dedicated to this argument.
Here is an overview of the argument in Section 8.1. Given $D\in\mathfrak{q}$, we lift $D$ to a $k$-algebra homomorphism $\partial D\colon A\to k[\delta]$ as in Proposition 7.3. Viewing $\partial D$ as a $k[\delta]$-point of $Q$ gives us $k[\delta]$-points $\operatorname{Ad}_{\partial D}\in\operatorname{GL}(k[\delta]\otimes_{k}\mathfrak{q})$ and $\operatorname{Ad}_{\partial D}^{-1}\in\operatorname{GL}(\mathfrak{q}\otimes_{k}k[\delta])$ of $\operatorname{GL}(\mathfrak{q})$.
Identify $k[\delta]\otimes_{k}\mathfrak{q}$ and $\mathfrak{q}\otimes_{k}k[\delta]$ with $\mathfrak{q}[\delta]/(\delta^{2})$. Then the second isomorphism in (7.1) states that for all $k$-derivations $E\in\mathfrak{q}$, the element of $\mathfrak{q}$ corresponding to $\operatorname{ad}_{D}(E)$ (resp. $\operatorname{ad}_{D}^{-1}(E)$) is precisely the coefficient of $\delta$ in $\operatorname{Ad}_{\partial D}(1\otimes E)$ (resp. $\operatorname{Ad}_{\partial D}^{-1}(E\otimes 1)$) in $\mathfrak{q}[\delta]/(\delta^{2})$. We show that this coefficient is the convolution $[D,E]$ (resp. $\{E,D\}$).
@@ 8.1. Convolution versus $\operatorname{ad}$
8.1.1.
In the following, fix arbitrary $k$-derivations $D,E\in\mathfrak{q}$. Our proofs of Propositions 8.2 and 9.6 will use the $k$-linear maps $\psi_{D},\varphi_{D}\colon A\to A$ defined by
(8.1) $\psi_{D}\coloneq(D\otimes\mathrm{id})\circ\nabla,\qquad\varphi_{D}\coloneq(\mathrm{id}\otimes D)\circ\nabla^{-1}.$
@@
Via the identifications $k[\delta]\otimes_{k}A\xrightarrow{\sim}A[\delta]$ and $A\otimes_{k}k[\delta]\xrightarrow{\sim}A[\delta]$, we obtain from (6.3) and (6.4) two automorphisms $L_{\partial D}^{\sharp},(L_{\partial D}^{-1})^{\sharp}$ of $A[\delta]$ defined by
$L_{\partial D}^{\sharp}(\delta^{n}f)=\delta^{n}(\partial D\otimes\mathrm{id}_{A})(\nabla f),\qquad(L_{\partial D}^{-1})^{\sharp}(\delta^{n}f)=\delta^{n}(\mathrm{id}_{A}\otimes\partial D)(\nabla^{-1}f)$
for all $f\in A$. The following gives alternative formulas for these automorphisms.
@@
$L_{\partial D}^{\sharp}(\delta^{n}f)=\delta^{n}(f+\delta\psi_{D}(f)),\qquad(L_{\partial D}^{-1})^{\sharp}(\delta^{n}f)=\delta^{n}(f+\delta\varphi_{D}(f)).$
@@
$=\delta^{n}\left(\sum\varepsilon(f_{(1)})f_{(2)}+\delta\sum D(f_{(1)})f_{(2)}\right)$
@@
where in the last equality we have used the co-fixing identity 4. The calculation for $(L_{\partial D}^{-1})^{\sharp}$ is similar and uses the co-inverse fixing identity 4. ∎
8.1.3.
Similarly, since $\partial D$ is a $k[\delta]$-point of $Q$ and $\operatorname{Ad}^{\pm 1}\colon Q\to\operatorname{GL}(\mathfrak{q})$ are morphisms of schemes, we obtain $k[\delta]$-points of $\operatorname{GL}(\mathfrak{q})$
$\operatorname{Ad}_{\partial D}\in\operatorname{GL}(k[\delta]\otimes_{k}\mathfrak{q}),\qquad\operatorname{Ad}_{\partial D}^{-1}\in\operatorname{GL}(\mathfrak{q}\otimes_{k}k[\delta]).$
Identify $k[\delta]\otimes_{k}A\xrightarrow{\sim}A[\delta]$ and $A\otimes_{k}k[\delta]\xrightarrow{\sim}A[\delta]$. Then the definition of the derivative shows for all $k$-derivations $E\in\mathfrak{q}$, we have
(8.3) $\operatorname{Ad}_{\partial D}(1\otimes E)=E\circ L_{\partial D}^{\sharp},\qquad\operatorname{Ad}_{\partial D}^{-1}(E\otimes 1)=E\circ(L_{\partial D}^{-1})^{\sharp}$
as maps $A[\delta]\to k[\delta]$. We use this to show the following.
@@
When considered as a map $\operatorname{ad}\colon\mathfrak{q}\to\mathfrak{gl}(\mathfrak{q})$, the adjoint representation of $\mathfrak{q}$ is precisely the map $D\mapsto[D,\cdot]$. That is, $\operatorname{ad}_{D}(E)=[D,E]$ for all $k$-derivations $D,E\in\mathfrak{q}$. Similarly, $\operatorname{ad}^{-1}_{D}(E)=\{E,D\}$ if $\operatorname{ad}^{-1}$ is considered as a map $\mathfrak{q}\to\mathfrak{gl}(\mathfrak{q})$.
In particular, $\mathfrak{q}$ is closed under $\cdot$$\cdot$] and $\{\cdot,\cdot\}$.
@@
We prove the claim for $\operatorname{ad}$ and $\cdot$$\cdot$]; the claim for $\operatorname{ad}^{-1}$ and $\{\cdot,\cdot\}$ is shown similarly.
@@
If $G$ is an affine algebraic group and $Q=\operatorname{Conj}(G)$ is its conjugation quandle scheme, then $(\mathfrak{q},[\cdot,\cdot])$ is the Lie algebra of $G$. To see this, note that $Q$ and $G$ have the same underlying scheme and identity $e$, so they also have the same tangent space $\mathfrak{q}$ at $e$. By the definition of $\operatorname{Conj}(G)$, the adjoint representations $\operatorname{Ad}$ of $Q$ and $\operatorname{ad}$ of $\mathfrak{q}$ are therefore the same as those of $G$ and its Lie algebra (see, for example, [milne]*Rem. 10.24). Hence, the claim follows from Proposition 8.2.444Alternatively, one can use Example 6.6 and the Hopf algebra axioms to directly (if tediously) verify that the Leibniz bracket $D,E]=(D\otimes E)\circ\nabla$ agrees with the usual Lie bracket $(D\otimes E-E\otimes D)\circ\Delta$ of $\mathfrak{q}=\operatorname{Lie}(G)$ (see [waterhouse]*Sec. 12.2); we leave the details to the reader. As a hint, first deduce the identity $\varepsilon\circ S=\varepsilon$ from the axiom $\mu\circ(S\otimes\mathrm{id})\circ\Delta=\eta\circ\varepsilon$ by post-composing with $\varepsilon$, and then deduce from these two equations that $D\circ S=-D$.